OpenStudy (anonymous):

Can somebody please show me how to do this? Thanks a bunch. Given a regular square pyramid with RS = 6 and PX = 4, find the following measure. XY = ?

ganeshie8 (ganeshie8):

there is no RS in the pic. must be a mistake check once

OpenStudy (anonymous):

okay i noticed that too and was wondering if that was just me or not. thanks!

ganeshie8 (ganeshie8):

lol !

ganeshie8 (ganeshie8):

one problem less in homework :)

OpenStudy (anonymous):

oh wait! lol good point tho, i just got the correct angle from my teacher. he said the tech. people put the wrong angle in the problem. one sec...

ganeshie8 (ganeshie8):

ok so no escape :(

OpenStudy (anonymous):

im sorry lol but do u think u can help me with it for maybe 1 minute? lolz. heres the correct angle.

OpenStudy (anonymous):

this angle match the problem better this time? :P

ganeshie8 (ganeshie8):

yeah this is good :) you want to find XY ?

ganeshie8 (ganeshie8):

XY is just half of RS XY = RS/2 = 6/2 = 3

OpenStudy (anonymous):

r u serious?? so... i just divide nothing else?

ganeshie8 (ganeshie8):

yes very much. for a regular square pyramid, altitude or PX goes thru center of square. so XY will be half of the side

OpenStudy (anonymous):

that makes sense. i hate it when they make problems look so complicated and they turn out so easy. GRR but thanks a bunch for ur help :)

ganeshie8 (ganeshie8):

haha np :) just call ur teacher and ask if its correct pic which she has sent this time. maybe she sent a wrong pic again lol

OpenStudy (anonymous):

geez i hope not :P